1 Purpose of Purchase

The Student Information Processing Board would like to obtain a 486 machine for the primary purpose of continued Linux development. In light of resnet, the SIPB would like to be able to provide a certain level of support for the Linux operating system to the user community, and having machines to develop on is necassary to provide this service.

2 Details of Purchase

The configuration we would like is as follows.

• 486DX

• 16M RAM

• 200M hard drive

• 15” Color monitor

• Ethernet Card

• 3 \frac{1}{2}” floppy

• Mouse

3 Cost

Given these spcifications, this system should cost approximately $4000 if purchased from the MCC, and somewhat less if purchased elsewhere.